OpenStudy (anonymous):

Solve this system algebraically. 9x + 2y = 5 y - 2x + 3 = 0 {(11/13, -17/13)} {(-11/13, -17/13)} {(-11/13, 17/13)} {(11/13, 17/13)}

OpenStudy (anonymous):

really confused

OpenStudy (acacia):

9x + 2y = 5 y - 2x + 3 = 0 Take the second equation, and solve it for y: y - 2x + 3 = 0 y = 2x - 3 Now plug in this value for y into the first equation: 9x + 2(2x-3) = 5 and solve for x: 9x + 4x - 6 = 5 13x = 11 x = (11/13) Now that you have x, plug it into either the first or second equation and solve for y. y = 2x - 3 y = 2(11/13) - 3 y = 22/13 - (39/13) y = -17/13 The answer to number 1 is: A

OpenStudy (anonymous):

thx helped lots

OpenStudy (acacia):

You're welcome